The probability of event ? is ?(?)=0.5 , and the probability of event ? is ?(?)=0.7 . Are ? and ? disjoint? no yes It is impossible to determine from the information given.
Answer #1

Ans: No

If A and B are disjoint, P(A \cap B) = 0

In that case, P(A U B) = P(A) + P(B) = 0.5 + 0.7 = 1.2

This is not possible. So, the value of P(A \cap B) must be at least 0.2, meaning that A and B are not disjoint.